OpenStudy (anonymous):

9/32 - 1/16 =? has to be a fraction

OpenStudy (uri):

Cross multiply ^

OpenStudy (anonymous):

\[\frac{ 9 }{ 32 }-\frac{ 1 }{ 16 }\]\[\frac{ 9 }{ 32 }-\frac{ 1*2 }{ 16*2 }\]\[\frac{ 9 }{ 32 }-\frac{ 2 }{ 32 }\]\[\frac{ 7 }{ 32 }\]

OpenStudy (anonymous):

the first line of the equation is the equation the was given

OpenStudy (anonymous):